We locked the PRMI, this time really on the sidebands, using the two REFL55 signals.
Here are the parameters: triggering on POP22_I in at 140, out at 20. No normalization. MICH gain -0.15, PRCL gain 0.1
It seems that the lock is not very stable. It seems likely to come from some large angular motion of one of the mirrors. We'll need to calibrate the optical lever signals to understand which one is moving too much.